Lux Enterprise is considering an investment project that generates a cash flow of $1,800,000 next year if the economy is favorable but generates only $900,000 if the economy is unfavorable. The probability of favorable economy is 55% and of unfavorable economy is 45%. The project will last only one year and be closed after that. The cost of investment is $1,300,000 and Lux Enterprise plans to finance the project with $350,000 of equity and $950,000 of debt. Assuming the discount rates of both equity and debt are 0%. What is the expected cash flow to Lux Enterprise s creditors if the enterprise invests in the project?
Answer #1

Cash Flow when Economy is Favorable = $1,800,000

Cash Flow when Economy is unfavorable = $900,000

Probability of Favorable Economy = 0.55

Probability of Unfavorable Economy = 0.45

Expected Cash Flow next year from Project = 0.55 * 1800000 + 0.45 * 900000

Expected Cash Flow next year from Project = 990,000 + 405,000

Expected Cash Flow next year from Project = $1,395,000

Cost of Project = $1,300,000

Debt Financing = $950,000

% Project is Debt Financed = 950000/1300000 = 73.07%

Cost of Debt and Equity = 0% (as per question)

Expected Cash Flow to creditors = 0.7307 * 1395000

Expected Cash Flow to creditors = $1,019,326.5